- Press Release
- Oct 2, 2022
NOAA GOES SATOPS Morning Report 13 November 2000
SATOPS MORNING REPORT – November 13, 2000
Some GOES instrument “speckling” (excitement of the instrument detectors
due to collisions with high energy particles) was observed last week that
were correlated to the coincident solar event. No additional performance
or health and safety concerns directly related to the event have been noted
on the POES, GOES or DMSP spacecraft.
No change in the status of the POES spacecraft.
NOAA-16 operations nominal. Reports from users indicate that the
APT signal (VTX transmitter turned on last Friday after the NOAA-14 conflict
period) is not as strong as expected. The VTX2 transmitter
turned on during rev 698 at 2155Z on Thursday, November 9. Engineers
will evaluate data today and discuss with NASA.
Engineers have developed and tested an on-board command MACRO to cycle
the NOAA-15 HIRS Filter Wheel as part of requested testing from NASA and
the instrument vendor, ITT. The objective of this command MACRO is
to toggle the Filter Wheel Off and On with a one second delay between the
two commands. This toggling will be performed multiple times during
the test period. Because this sequence impacts the HIRS Filter Wheel
only, user products should not be affected. The MACRO will be loaded
at ~17:09Z and the filter wheel will be cycled off for one second at approximately
18:51Z, 2211Z, 2333Z on November 13 and at 1648Z, 1828Z and 2008Z on November
A NOAA-15 solar array slew from -45 degrees to +53 degrees is planned
for Tuesday, November 14 as part of the continuing AVHRR (attempted) restoration
efforts. This commanding will occur on Rev 13029 / F, at 1827Z.
NOAA-14 rev 30214 / F at 1643Z on November 9: G3A noisy from spacecraft.
Received 71% of the data. Data recovery not possible due to recorder
NOAA-14 rev 30224 / W at 0952Z on November 10: L2B noisy from
spacecraft. Received 96.3% of the data. Data recovery not possible
due to recorder management.
NOAA-14 rev 30243 / F at 1800Z on November 11: G1B noisy from spacecraft.
Received 91% of the data. Data recovery not possible due to recorder
NOAA-14 rev 30254 / F at 1245Z on November 12: G1A noisy from
spacecraft. Received 96% of the data. Data recovery not possible
due to recorder management. This playback resembled three other noisy/short
playbacks in the past few days. All playbacks were GAC data from
N-14 on the LSB at the FCDAS and were the last playback of the pass.
These events will be reviewed this week.
NOAA-15 rev 12997 / W at 1235Z on November 12: G1A noisy from spacecraft.
Received 92% of the data. Data recovery not possible due to recorder
POES operations were nominal over the past 96 hours.
No change in the status of the DMSP spacecraft.
DMSP operations were nominal over the past 96 hours.
No change in the status of the GOES spacecraft.
At 0117Z, 0757Z and 2035Z on Saturday, November 11, GOES-10 frame breaks
were observed and attributed to loss of signal sync.
At 2333Z on Saturday, November 11, and at 0805Z on Sunday, November
12, GOES-8 frame breaks were observed and attributed to RFI at CDA.
At 0206Z on Monday, November 13, a GOES-8 frame break was observed and
attributed to loss of signal sync.
Throughout the day on November 9, the PM operator reported “spotting”
in the imagery (both GOES-8 and GOES-10); attributed to the solar storm
GOES-9 and GOES-11 remain in the Z-Axis Precession (ZAP) storage mode
at ~105 deg W.
GOES operations were nominal over the past 96 hours.
Questions or comments to email@example.com